### v3

@@ Abstract.
2020 Mathematics Subject Classification:
37B10, 68R15
The first author was supported by a FNRS fellowship.
1. Introduction
Dendric shifts are defined in terms of extension graphs that describe the left and right extensions of their factors. Extension graphs are bipartite graphs that can roughly be described as follows: if $u$ is a word in the language $\mathcal{L}(X)$ of the shift space $X$, one puts an edge between the left and right copies of letters $a$ and $b$ such that $aub$ is in $\mathcal{L}(X)$. A shift space is then said to be dendric if the extension graph of every word of its language is a tree. These shift spaces were initially defined through their languages under the name of tree sets [5] and were studied in a series of papers. They generalize classical families of shift spaces such as Sturmian shifts [30], Arnoux-Rauzy shifts [2], codings of regular interval exchange transformations (IET) [31, 1] , and shift spaces arising from the application of the Cassaigne multidimensional continued fraction algorithm (MCF) [13].
Minimal dendric shifts exhibit striking combinatorial [8, 10], algebraic [5, 7] , and ergodic properties [4]. For instance, they have factor complexity #$\#(\mathcal{L}(X)\cap\mathcal{A}^{n})=(\#\mathcal{A}-1)n+1$ [5] and topological rank #$\#\mathcal{A}$ [4], where $\mathcal{A}$ is the alphabet of the shift space. They also fall into the class of eventually dendric shift spaces, which are exactly those satisfying the regular bispecial condition [18]. This implies that the number of their ergodic measures is at most #$\#\mathcal{A}/2$ [16]. An important property for our work is that the derived shift of a minimal dendric shift is again a minimal dendric shift on the same alphabet, where derivation is here understood as derivation by return words. This allows to give $\mathfrak{S}$-adic representations of such shift spaces [26], i.e., to define a set $\mathfrak{S}$ of endomorphisms of the free monoid $\mathcal{A}^{*}$ and a sequence ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}\in\mathfrak{S}^{\mathbb{N}}$, called an $\mathfrak{S}$-adic representation, such that
@@ 1. Introduction
$\mathfrak{S}$-adic representations are a classical tool that allows to study several properties of shift spaces such as the factor complexity [24, 19], the number of ergodic measures [9, 3, 12], the dimension group and topological rank [4] or yet the automorphism group [25]. In the case of minimal dendric shifts, the involved endomorphisms are particular tame automorphisms of the free group generated by the alphabet [8, 10]. This in particular allows to prove that minimal dendric shifts have topological rank equal to the cardinality of the alphabet and that ergodic measures are completely determined by the measures of the letter cylinders [4, 12].
In the case of the ternary alphabet $\{1,2,3\}$, we were able to give an $\mathfrak{S}$-adic characterization of minimal dendric shifts [29]. We explicitly define an infinite set $\mathfrak{S}_{3}$ of positive automorphisms of the free group $F_{3}$ and a labeled directed graph $\mathcal{G}$ with two vertices and labels in $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$ (where $\Sigma_{3}$ is the symmetric group of $\{1,2,3\}$) such that a shift space over $\{1,2,3\}$ is a minimal dendric shift if and only if it has a primitive $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$-adic representation labelling an infinite path in $\mathcal{G}$. We were then able to localize in $\mathcal{G}$ classical families of dendric shifts such as codings of regular interval exchanges or Arnoux-Rauzy shifts.
@@
The paper is organized as follows. We start by giving, in Section 2, the basic definitions for the study of shift spaces. We introduce the notion of extension graph of a word and of a dendric shift. In Section 3, we recall and extend some results of [29] about the stability of dendricity when taking the image under a return morphism.
@@
2. Definitions
2.1. Words, languages and shift spaces
@@
2.2. Extension graphs and dendric shifts
@@
Proposition 2 (Cassaigne and Nicolas [14]).
@@
(1) $\operatorname{Card}(E(w))-\operatorname{Card}(E^{L}(w))-\operatorname{Card}(E^{R}(w))+1=0,$
@@
Figure 1. The extension graphs of $\varepsilon$ (on the left), 00 (in the center) and $1$ (on the right) are trees.
3. Extension graphs in morphic images
@@
    (2) $E_{Y}(u)=\{(a^{\prime},b^{\prime})\in\mathcal{B}\times\mathcal{B}\mid\exists(a,b)\in E_{X,s,p}(v):\sigma(a)\in\mathcal{B}^{*}a^{\prime}s\wedge\sigma(b)w\in pb^{\prime}\mathcal{B}^{*}\}.$
@@
For the second case, the proof is similar to the proof of [29, Proposition 4.1]. ∎
@@
The following result is [29, Proposition 4.8].
@@
Proposition 17 (Dolce and Perrin [18]).
@@
4. Other graphs characterizing dendricity
@@
Figure 2. Graphs $G_{n}^{L}(X)$ for $n\in\{0,\dots,3\}$.
@@
The result is false if we remove the acyclic hypothesis. Indeed, if $X$ is such that $\mathcal{L}_{3}(X)=\{001,010,011,100,110,111\}$, then the graph $\mathcal{E}_{X}(0)$ is not connected. However, $G^{L}_{2}(X)$ is and, indeed, the graph $\mathcal{E}_{X}(\varepsilon)$ contains a cycle (see Figure 3) thus this does not contradict the previous result.
@@
Figure 3. The acyclicity of $\mathcal{E}_{X}(v)$ is necessary in Proposition 27.
@@
5. Stabilization and eventually dendric shifts
@@
If $C=\{0,1,2,3\}$, $C_{1}=\{0,1\}$, $C_{2}=\{1,2,3\}$, $C_{3}=\{0,3\}$, the multi-clique $G(\{C_{1},C_{2},C_{3}\})$ is represented in Figure 4. It is not acyclic for the coloring because of the cycle $0,1,3,0$.
@@
Figure 4. Multi-clique $G(\{\{0,1\},\{1,2,3\},\{0,3\}\})$
@@
The Chacon ternary shift space over $\{0,1,2\}$ is generated by the morphism $\sigma:0\mapsto 0012,1\mapsto 12,2\mapsto 012$. It is of complexity $12n+$ 1 [28] thus, by Proposition 2, for each length $n$, there is either a unique left special word with three extensions or two left special words with two extensions each. The graph $G^{L}_{n}(X)$ then contains three or two edges respectively. Since the graph $G^{L}_{1}(X)$ only contains two edges between 00 and $2$ (labeled by 00 and $1$), using Remark 22, it will be the case of $G^{L}_{n}(X)$ for all $n\geq 1$. We can similarly show that the graphs $G^{R}_{n}(X)$ correspond to the same colored graph for all $n\geq 1$. However, the Chacon shift space is not eventually dendric [18].
@@ 6. Action of morphisms on $G^{L}(X)$ and $G^{R}(X)$
Figure 5. Graphs $G=G(\{\{0,1\},\{1,2,3\}\})$ (on the left) and $\beta^{L}(G)$ (on the right)
@@
7. $S$-adic characterization of minimal dendric shifts
@@ 7.1. $S$-adic representations
We say that the sequence ${\boldsymbol{\sigma}}$ is primitive if, for any $n$, there exists $N>n$ such that for all $(a,b)\in\mathcal{A}_{n}\times\mathcal{A}_{N}$, $a$ occurs in $\sigma_{[n,N)}(b)$. Observe that if ${\boldsymbol{\sigma}}$ is primitive, then $\min_{a\in\mathcal{A}_{n}}|\sigma_{[0,n)}(a)|$ goes to infinity when $n$ increases, $\mathcal{L}(X_{{\boldsymbol{\sigma}}}^{(n)})=\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$, and $X_{{\boldsymbol{\sigma}}}^{(n)}$ is a minimal shift space (see for instance [21, Lemma 7]).
@@
We let $\mathcal{R}_{X}(w)$ denote the set of return words to $w$ in $X$ and we omit the subscript $X$ whenever it is clear from the context. The shift space $X$ being minimal, $\mathcal{R}(w)$ is finite (see [20]) thus we write $R_{X}(w)=\{1,\dots,\operatorname{Card}(\mathcal{R}_{X}(w))\}$. A morphism $\sigma:{R(w)}^{*}\to\mathcal{A}^{*}$ is a coding morphism associated with $w$ if $\sigma(R(w))=\mathcal{R}(w)$. It is trivially a return morphism for $w$.
@@
7.2. Characterization via two graphs
Similarly to what was done in the case of a ternary alphabet in [29], using the results of Section 6, we can give an $S$-adic characterization of the minimal dendric shifts on an alphabet $\mathcal{A}$. This characterization uses two graphs corresponding to the left and the right extensions respectively.
@@ 7.2. Characterization via two graphs
Now assume that ${\boldsymbol{\sigma}}$ is primitive and labels a path $\left(G^{L}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{L}(\mathfrak{S})$ and a path $\left(G^{R}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{R}(\mathfrak{S})$. The shift $X_{\boldsymbol{\sigma}}$ is minimal by primitiveness of ${\boldsymbol{\sigma}}$ [21]. Let $u\in\mathcal{L}(X_{\boldsymbol{\sigma}})$ be a factor. If $u$ is a $\sigma_{0}$-initial factor, then it is dendric since $\sigma_{0}$ is dendric. Otherwise, iterating Proposition 8, there exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$. Let $Y$ be an Arnoux-Rauzy shift and let $Z=\sigma_{k}\cdot Y$. Since $v$ is $\sigma_{k}$-initial, $\mathcal{E}_{Z}(v)=\mathcal{E}_{X_{\boldsymbol{\sigma}}^{(k)}}(v)$. As $Y$ is an Arnoux-Rauzy shift, the graph $G^{L}(Y)$ (resp., $G^{R}(Y)$) is the complete (simple and monochromatic) graph on $\mathcal{A}$, thus $G^{L}_{k+1}$ (resp., $G^{R}_{k+1}$) is a covering subgraph. By Lemma 51 applied $k+1$ times, we obtain that $\sigma_{0}\dots\sigma_{k-1}\cdot Z$ is dendric thus, starting with the extension graph $\mathcal{E}_{Z}(v)$, $v$ only has dendric extended images under $\sigma_{0}\dots\sigma_{k-1}$ and $u$ is dendric. ∎
@@
In the case of a ternary alphabet, this allows to replace a graph with 9 vertices by a graph with 2 vertices, as done in [29]. For an alphabet of size 4, we obtain a graph with 14 vertices instead of 256.
7.3. An example: $\mathfrak{S}$-adic characterization of minimal dendric shifts with one right special factor of each length
@@
Figure 6. Possible extension graphs of the empty word in a minimal dendric shift over $\{0,1,2,3\}$ and their associated return morphisms.
@@
Figure 7. Graphs $G_{1}$ on the left and $G_{2}$ on the right
@@
Figure 8. A shift space over $\{0,1,2,3\}$ is minimal, dendric and has exactly one right special factor of each length if and only if it has a primitive $\mathfrak{S}_{1\text{R}}$-adic representation labeling an infinite path in this graph.
@@
7.4. Characterization of eventually dendric shifts

### v4

@@ Abstract
1 Introduction
Dendric shifts are defined in terms of extension graphs that describe the left and right extensions of their factors. Extension graphs are bipartite graphs that can roughly be described as follows: if $u$ is a word in the language $\mathcal{L}(X)$ of the shift space $X$, one puts an edge between the left and right copies of letters $a$ and $b$ such that $aub$ is in $\mathcal{L}(X)$. A shift space is then said to be dendric if the extension graph of every word of its language is a tree. These shift spaces were initially defined through their languages under the name of tree sets ((Berthé et al., 2015a)) and were studied in a series of papers. They generalize classical families of shift spaces such as Sturmian shifts ((Morse and Hedlund, 1940)), Arnoux-Rauzy shifts ((Arnoux and Rauzy, 1991)), codings of regular interval exchange transformations (IET) ((Oseledets, 1966; Arnold, 1963)) , and shift spaces arising from the application of the Cassaigne multidimensional continued fraction algorithm (MCF) ((Cassaigne et al., 2022)).
Minimal dendric shifts exhibit striking combinatorial ((Berthé et al., 2015c, 2018)), algebraic ((Berthé et al., 2015a, d)) , and ergodic properties ((Berthé et al., 2021)). For instance, they have factor complexity #$\#(\mathcal{L}(X)\cap\mathcal{A}^{n})=(\#\mathcal{A}-1)n+1$ ((Berthé et al., 2015a)) and topological rank #$\#\mathcal{A}$ ((Berthé et al., 2021)), where $\mathcal{A}$ is the alphabet of the shift space. They also fall into the class of eventually dendric shift spaces, which are exactly those satisfying the regular bispecial condition ((Dolce and Perrin, 2021)). This implies that the number of their ergodic measures is at most #$\#\mathcal{A}/2$ ((Damron and Fickenscher, 2022)). An important property for our work is that the derived shift of a minimal dendric shift is again a minimal dendric shift on the same alphabet, where derivation is here understood as derivation by return words. This allows to give $\mathfrak{S}$-adic representations of such shift spaces ((Ferenczi, 1996)), i.e., to define a set $\mathfrak{S}$ of endomorphisms of the free monoid $\mathcal{A}^{*}$ and a sequence ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}\in\mathfrak{S}^{\mathbb{N}}$, called an $\mathfrak{S}$-adic representation, such that
@@ 1 Introduction
$\mathfrak{S}$-adic representations are a classical tool that allows to study several properties of shift spaces such as the factor complexity ((Durand et al., 2013; Donoso et al., 2021)), the number of ergodic measures ((Berthé and Delecroix, 2014; Bédaride et al., 2021, 2020)), the dimension group and topological rank ((Berthé et al., 2021)) or yet the automorphism group ((Espinoza and Maass, 2022)). In the case of minimal dendric shifts, the involved endomorphisms are particular tame automorphisms of the free group generated by the alphabet ((Berthé et al., 2015c, 2018)). This in particular allows to prove that minimal dendric shifts have topological rank equal to the cardinality of the alphabet and that ergodic measures are completely determined by the measures of the letter cylinders ((Berthé et al., 2021; Bédaride et al., 2020)).
In the case of the ternary alphabet $\{1,2,3\}$, we were able to give an $\mathfrak{S}$-adic characterization of minimal dendric shifts ((Gheeraert et al., 2022)). We explicitly define an infinite set $\mathfrak{S}_{3}$ of positive automorphisms of the free group $F_{3}$ and a labeled directed graph $\mathcal{G}$ with two vertices and labels in $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$ (where $\Sigma_{3}$ is the symmetric group of $\{1,2,3\}$) such that a shift space over $\{1,2,3\}$ is a minimal dendric shift if and only if it has a primitive $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$-adic representation labelling an infinite path in $\mathcal{G}$. We were then able to localize in $\mathcal{G}$ classical families of dendric shifts such as codings of regular interval exchanges or Arnoux-Rauzy shifts.
@@
The paper is organized as follows. We start by giving, in Section 2, the basic definitions for the study of shift spaces. We introduce the notion of extension graph of a word and of a dendric shift. In Section 3, we recall and extend some results of ((Gheeraert et al., 2022)) about the stability of dendricity when taking the image under a return morphism.
@@
2 Definitions
2.1 Words, languages and shift spaces
@@
2.2 Extension graphs and dendric shifts
@@
Proposition 2 (Cassaigne and Nicolas ((Cassaigne and Nicolas, 2010))).
@@
$\operatorname{Card}(E(w))-\operatorname{Card}(E^{L}(w))-\operatorname{Card}(E^{R}(w))+1=0,$ (1)
@@
Figure 1: The extension graphs of $\varepsilon$ (on the left), 00 (in the center) and $1$ (on the right) are trees.
3 Extension graphs in morphic images
@@
    $E_{Y}(u)=\{(a^{\prime},b^{\prime})\in\mathcal{B}\times\mathcal{B}\mid\exists(a,b)\in E_{X,s,p}(v):\sigma(a)\in\mathcal{B}^{*}a^{\prime}s\wedge\sigma(b)w\in pb^{\prime}\mathcal{B}^{*}\}.$ (2)
@@
For the second case, the proof is similar to the proof of ((Gheeraert et al., 2022, Proposition 4.1)). ∎
@@
The following result is ((Gheeraert et al., 2022, Proposition 4.8)).
@@
Proposition 17 (Dolce and Perrin ((Dolce and Perrin, 2021))).
@@
4 Other graphs characterizing dendricity
@@
Figure 2: Graphs $G_{n}^{L}(X)$ for $n\in\{0,\dots,3\}$.
@@
The result is false if we remove the acyclic hypothesis. Indeed, if a shift space $X$ is such that $\mathcal{L}_{3}(X)=\{001,010,011,100,110,111\}$, then the graph $\mathcal{E}_{X}(0)$ is not connected. However, $G^{L}_{2}(X)$ is and, indeed, the graph $\mathcal{E}_{X}(\varepsilon)$ contains a cycle (see Figure 3) thus this does not contradict the previous result.
@@
Figure 3: The acyclicity of $\mathcal{E}_{X}(v)$ is necessary in Proposition 27.
@@
5 Stabilization and eventually dendric shifts
@@
If $C=\{0,1,2,3\}$, $C_{1}=\{0,1\}$, $C_{2}=\{1,2,3\}$ and $C_{3}=\{0,3\}$, the multi-clique $G(\{C_{1},C_{2},C_{3}\})$ is represented in Figure 4. It is not acyclic for the coloring because of the cycle $0,1,3,0$.
@@
Figure 4: Multi-clique $G(\{\{0,1\},\{1,2,3\},\{0,3\}\})$
@@
The Chacon ternary shift space over $\{0,1,2\}$ is generated by the morphism $\sigma:0\mapsto 0012,1\mapsto 12,2\mapsto 012$. It is of complexity $12n+$ 1 ((Fogg, 2002)) thus, by Proposition 2, for each length $n$, there is either a unique left special word with three extensions or two left special words with two extensions each. The graph $G^{L}_{n}(X)$ then contains three or two edges respectively. Since the graph $G^{L}_{1}(X)$ only contains two edges between 00 and $2$ (labeled by 00 and $1$), using Remark 22, it will be the case of $G^{L}_{n}(X)$ for all $n\geq 1$. We can similarly show that the graphs $G^{R}_{n}(X)$ correspond to the same colored graph for all $n\geq 1$. However, the Chacon shift space is not eventually dendric ((Dolce and Perrin, 2021)).
@@ 6 Action of morphisms on $G^{L}(X)$ and $G^{R}(X)$
Figure 5: Graphs $G=G(\{\{0,1\},\{1,2,3\}\})$ (on the left) and $\beta^{L}(G)$ (on the right)
@@
7 $S$-adic characterization of minimal dendric shifts
@@ 7.1 $S$-adic representations
We say that the sequence ${\boldsymbol{\sigma}}$ is primitive if, for any $n$, there exists $N>n$ such that for all $(a,b)\in\mathcal{A}_{n}\times\mathcal{A}_{N}$, $a$ occurs in $\sigma_{[n,N)}(b)$. Observe that if ${\boldsymbol{\sigma}}$ is primitive, then $\min_{a\in\mathcal{A}_{n}}|\sigma_{[0,n)}(a)|$ goes to infinity when $n$ increases, $\mathcal{L}(X_{{\boldsymbol{\sigma}}}^{(n)})=\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$, and $X_{{\boldsymbol{\sigma}}}^{(n)}$ is a minimal shift space (see for instance ((Durand, 2000, Lemma 7))).
@@
We let $\mathcal{R}_{X}(w)$ denote the set of return words to $w$ in $X$ and we omit the subscript $X$ whenever it is clear from the context. The shift space $X$ being minimal, $\mathcal{R}(w)$ is finite (see ((Durand, 1998))) thus we write $R_{X}(w)=\{1,\dots,\operatorname{Card}(\mathcal{R}_{X}(w))\}$. A morphism $\sigma:{R(w)}^{*}\to\mathcal{A}^{*}$ is a coding morphism associated with $w$ if $\sigma(R(w))=\mathcal{R}(w)$. It is trivially a return morphism for $w$.
@@
7.2 Characterization via two graphs
Similarly to what was done in the case of a ternary alphabet in ((Gheeraert et al., 2022)), using the results of Section 6, we can give an $S$-adic characterization of the minimal dendric shifts on an alphabet $\mathcal{A}$. This characterization uses two graphs corresponding to the left and the right extensions respectively.
@@ 7.2 Characterization via two graphs
Now assume that ${\boldsymbol{\sigma}}$ is primitive and labels a path $\left(G^{L}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{L}(\mathfrak{S})$ and a path $\left(G^{R}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{R}(\mathfrak{S})$. The shift $X_{\boldsymbol{\sigma}}$ is minimal by primitiveness of ${\boldsymbol{\sigma}}$ ((Durand, 2000)). Let $u\in\mathcal{L}(X_{\boldsymbol{\sigma}})$ be a factor. If $u$ is a $\sigma_{0}$-initial factor, then it is dendric since $\sigma_{0}$ is dendric. Otherwise, iterating Proposition 8, there exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$. Let $Y$ be an Arnoux-Rauzy shift and let $Z=\sigma_{k}\cdot Y$. Since $v$ is $\sigma_{k}$-initial, $\mathcal{E}_{Z}(v)=\mathcal{E}_{X_{\boldsymbol{\sigma}}^{(k)}}(v)$. As $Y$ is an Arnoux-Rauzy shift, the graph $G^{L}(Y)$ (resp., $G^{R}(Y)$) is the complete (simple and monochromatic) graph on $\mathcal{A}$, thus $G^{L}_{k+1}$ (resp., $G^{R}_{k+1}$) is a covering subgraph. By Lemma 51 applied $k+1$ times, we obtain that $\sigma_{0}\dots\sigma_{k-1}\cdot Z$ is dendric thus, starting with the extension graph $\mathcal{E}_{Z}(v)$, $v$ only has dendric extended images under $\sigma_{0}\dots\sigma_{k-1}$ and $u$ is dendric. ∎
@@
In the case of a ternary alphabet, this allows to replace a graph with 9 vertices by a graph with 2 vertices, as done in ((Gheeraert et al., 2022)). For an alphabet of size 4, we obtain a graph with 14 vertices instead of 256.
7.3 An example: $\mathfrak{S}$-adic characterization of minimal dendric shifts with one right special factor of each length
@@
Figure 6: Possible extension graphs of the empty word in a minimal dendric shift over $\{0,1,2,3\}$ and their associated return morphisms.
@@
Figure 7: Graphs $G_{1}$ on the left and $G_{2}$ on the right
@@
Figure 8: A shift space over $\{0,1,2,3\}$ is minimal, dendric and has exactly one right special factor of each length if and only if it has a primitive $\mathfrak{S}_{1\text{R}}$-adic representation labeling an infinite path in this graph.
@@
7.4 Characterization of eventually dendric shifts